Water is flowing continuously from a tap having an internal diameter 8*10-3m. The water velocity as it leaves the tap is 0.4 ms-1. The diameter of the water stream at a distance 2*10-1m below the tap is close to :
(a) 7.5*10-3m
(b) 9.6*10-3m
(c) 3.6*10-3m
(d) 5.4*10-3m

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
(c). From Bernoulli’s theorem,
P0 + 1/2 pv21pgh = p0 + 1/2 pv22 + 0
V2 = √v21 + 2gh = √0.16 + 2 * 10 *0.2 = 2.03 m/s
From equation of continuity
A2v2 = A1v1
πD22/4 * v2 = πD21/4v1
\RightarrowD1D2 √v1/v2 = 3.55 * 10-3 m